find three consecutive odd integers such that the sum of the first and twice the second is 6 more than the third.
Answer by rfadrogane(214)

find three consecutive odd integers such that the sum of the first and twice the second is 6 more than the third.
Sol'n:
let: x - the 1st integer
x+2 - the 2nd integer
x+4 - the 3rd integer
x+2(x+2)= (x+4)+6
3x+4=x+10
2x=10-4
2x=6, divide by 2 on both side
x=3
x+2=5
x+4=7
thus,the add integer are the ff.
3,5&7 ----answer